Find the real functions and eigenvalues of the problem

{%O y + 6y + (5 + X)y = 0 pour 0 < x < 2, y(0) = 0 et y(2) = 0.
We ask for a complete solution where each step is justified
Answer #1

y+64 + (54) 4.0 ; y los= 0, y)=0 The characteristic equation of the above differential equation es m² tomt 5t-o - 6 + 136-2; NEN. η. Case al -32 and m₂ = -3-k fc6-2 k) = 0 solution of the differential est. in y (x)= sin Hence the eigen function ofon solving equation and ③ = 4 = 0 . In this case we also get trivial sol. ie. y(x) = Hence the only case case is possible wh